HUBBLE SPACE TELESCOPE DAILY REPORT #1494 PERIOD COVERED: 0000Z (UTC) 10/11/95 - 0000Z (UTC) 10/12/95 Daily Status Report as of 285/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6225 (Resolving Dust Around Shell Stars) The WF/PC-2 was used to make deep image observations of A-class = shell star HR-2174 looking for protoplanetary disks, like that around Beta = Pic. The observations were executed as scheduled, and no anomalies were = noted. 1.2 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flats) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.3 Completed WF/PC-2 6113 (Searching for Low Mass Stars: The Mass = Function ) The WF/PC-2 was used to make observations of globular cluster = NGC-6254 to search for very low mass (VLM) stars. The intent is to derive = the luminosity and mass functions for VLM stars. The observations were = completed as planned, and no problems were reported. 1.4 Completed Two Sets of WF/PC-2 6119 (Snapshot Survey for Companions = of Planetary Nebula Nuclei) The WF/PC-2 was used to make snapshot survey observations of = planetary nebula EGB6 to study its nuclei and to look for possible = companions. The observations were executed as scheduled, and no anomalies = were noted. 1.5 Completed GHRS/2 5961 (The D/H Ratio in Double QSO Absorption = System Towards Q0454+220) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the Lyman Alpha line toward QSO Q0454-220 = in order to study the Deuterium to Hydrogen ratio. A wavelength = calibration image was also taken. The observations were completed as = planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 03 Successful : 03 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: No problems were reported. All observatiory subsystems performinf = nominally. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-